Please help.
Venture capital. Henry invested $12,000 in a new restaurant.When the restaurant was sold two years later, he recieved $27,000. Find his average annual return by solving the equation 12,000(1+r)^2=27,000.

Answer by checkley75(3666)

12,000(1+r)^2=27,000.
12,000(1+2r+r^2)=27,000
12,000+24,000r+12,000r^2=27,000
12,000r^2+24,000r+12,000-27,000=0
12,000r^2+24,000r-15,000=0
3,000(4r^2+8r-5)=0
3,000(2r+5)(2r-1)=0
2R-1=0
2R=1
R=1/2=.5 OR 50%.
pROOF:
12,000(1,5)^2=27,000
12,000*2.25=27,000
27,000=27,000